OpenStudy (epicteatime):

227 229 271 317 349 You can use all the types of methods to try and get the answer, but you MUST use 1997, and keep it in that order no matter what. Please help, too hard 4 me...

OpenStudy (epicteatime):

@whpalmer4

OpenStudy (epicteatime):

@ParthKohli

OpenStudy (epicteatime):

@Luigi0210

OpenStudy (epicteatime):

@dan815

OpenStudy (epicteatime):

@UnkleRhaukus

OpenStudy (unklerhaukus):

What is the question?

OpenStudy (epicteatime):

You have to make those numbers by using 1997

OpenStudy (epicteatime):

you can't rearrange them, but you can use anything you want to make it the number. You can use powers, square roots, etc.

OpenStudy (epicteatime):

I can't get it at all, so I was wondering if anyone else could help me with this xD

imqwerty (imqwerty):

interesting :)

OpenStudy (epicteatime):

Can't do it xD

OpenStudy (epicteatime):

Closest I got was 220

imqwerty (imqwerty):

can we rearrange the number 1,9,9,7?

OpenStudy (epicteatime):

nope xD

OpenStudy (epicteatime):

it has to stay in that specific order

OpenStudy (unklerhaukus):

(1 + 9) x 9 x √7 = 238.1

OpenStudy (epicteatime):

SO CLOSE

OpenStudy (unklerhaukus):

this is hard, how did you get 220?

OpenStudy (epicteatime):

(19+9)+7^2

OpenStudy (unklerhaukus):

(1 + √9) x 9 x 7 = 252

OpenStudy (unklerhaukus):

(19 x 9) + 7^2 = 220

OpenStudy (epicteatime):

yep, thats what I did xD

OpenStudy (epicteatime):

closest thing I got xD

OpenStudy (unklerhaukus):

(1 x 9 x 9) + √7 = 214.3

OpenStudy (epicteatime):

:o

OpenStudy (unklerhaukus):

\[\sqrt{\sqrt{1\times 9^9\times7}}=228.2\]

OpenStudy (epicteatime):

wow, thats really close /:

OpenStudy (epicteatime):

So annoying... this stupid assignment.

OpenStudy (anonymous):

can you use powers of numbers other than 1,9,9,7?

OpenStudy (epicteatime):

if you mean, can you square it or "double it" yes you can

OpenStudy (epicteatime):

but 1997 has to stay in place/ keep the same structure.

OpenStudy (anonymous):

so you can use 1997 as many times as you want but it keeps that structure, like 1*9*9*7*1*9*9....etc?

OpenStudy (epicteatime):

oh... no thats not usable xD

OpenStudy (anonymous):

so you can only use 1997 once then?

OpenStudy (epicteatime):

mhmm

OpenStudy (anonymous):

(1 x 9 x 9) + √7 = 214.3

OpenStudy (anonymous):

but wouldn't it be defeating the purpose if you use powers of 3 then? even tho you are multiplying it 3 times... your using a number not in that set..

OpenStudy (anonymous):

probs over thinking it.

OpenStudy (epicteatime):

uhhh xD

OpenStudy (epicteatime):

I dunno xD Its hard just trying to do it the legit way xD

OpenStudy (jadedry):

227 is a prime number, so I suppose the manipulation of the numbers you have should reflect that... I'm struggling quite a bit as well, but maybe you should work backwards? add 1, 9 or 7 to your target number and try to factorize it to get hints on how to create it..

OpenStudy (epicteatime):

:o thats a good idea, i'll try that now xD

OpenStudy (epicteatime):

I just got 226 :P

OpenStudy (epicteatime):

gg no re jade is a genius

imqwerty (imqwerty):

can we do something like this- \[1 \times \sqrt{9},\sqrt[\infty]{9},7 =317\]?

OpenStudy (epicteatime):

yep

imqwerty (imqwerty):

so 317 is done :)

OpenStudy (epicteatime):

:o thank you xD

OpenStudy (epicteatime):

But I don't get how it becomes 317 xD

imqwerty (imqwerty):

\[1 \times \sqrt{9}=1 \times 3=3\]\[\sqrt[\infty]{9}=1\]\[7=7\] joining all :) 317

OpenStudy (epicteatime):

:o

OpenStudy (epicteatime):

thank you for your explanation xD

imqwerty (imqwerty):

np (:

OpenStudy (epicteatime):

and thank you jade, i'll make sure to use your methods :D

OpenStudy (jadedry):

@imqwerty You can join the numbers? don't you have to add or multiply?

OpenStudy (epicteatime):

OH

imqwerty (imqwerty):

idk? @Epicteatime said yes so i did it :)

OpenStudy (epicteatime):

Yeah, you have to add them or multiply them to get the actual number, sorry Jade I didnt see how you got it xD

OpenStudy (epicteatime):

not jade, imqwerty

OpenStudy (anonymous):

1×9√,9√∞,7=317

OpenStudy (jadedry):

ah, so no joining then, that would have made it easier. :c

OpenStudy (anonymous):

2^2 • 3 • 19 = 228 - (227 +1) 2 • 113 = 226 = (227-1) 22 • 5 • 11 = 220 = (227 - 7) 2 • 32 • 13 = 234 = (227 +7) 22 • 59 = 236 = (227 +9) 2 • 109 = 218 = (227 - 9)

OpenStudy (epicteatime):

yeah, it would have

OpenStudy (epicteatime):

lauren, thats not how you do it xD Thanks for your effort though :D

OpenStudy (jadedry):

I'm sure this is actually a simple question at it's heart, I doubt we need fancy, doctorate level thesis' on how this could work. The numbers we have are 1 , 9 , 9 , 7 , they have to be dealt with in that order, but they can be subtracted and added to each other, as well as divided and multiplied. We can use any number we wish when "doubling" or calculating to the power, right @epicteatime ?

OpenStudy (jadedry):

@Epicteatime I got ONE answer. \[1 * 9^{2.5} - 9 - 7 = 227 \]

OpenStudy (epicteatime):

sorry, I went to eat dinner xD Im back now tho :D

OpenStudy (epicteatime):

and thank you for answering one xD

OpenStudy (epicteatime):

sadly there's still like, 4 :P

OpenStudy (jadedry):

@Epicteatime Hope you enjoyed dinner haha, I know, but once one goes down, surely the rest should be easier? XD

OpenStudy (epicteatime):

maybe, maybe

OpenStudy (epicteatime):

xD

OpenStudy (epicteatime):

which one are you doing jade?

OpenStudy (jadedry):

I'm trying to do the second one now, it's only a difference of 2 so...

OpenStudy (epicteatime):

alright, i'll do 271 then :D

OpenStudy (epicteatime):

holy crap...

OpenStudy (epicteatime):

I Just got 271

OpenStudy (epicteatime):

but I still need to do something with 7 xD

OpenStudy (jadedry):

So, I need to subtract 14 from \[9^{2.5}\] in order to get 229. How can I make 14? --- Also, congrats! I'm sure you can turn the 7 into a 0 somehow...

OpenStudy (firekat97):

so which ones do you have left? haha

OpenStudy (epicteatime):

229 271 317 349

OpenStudy (epicteatime):

those :3

OpenStudy (firekat97):

ahh okiee

OpenStudy (epicteatime):

yush :3

OpenStudy (epicteatime):

I GOT 271!!!!

OpenStudy (epicteatime):

oh wait...

OpenStudy (epicteatime):

1997-99-1/7

OpenStudy (epicteatime):

uhh

OpenStudy (epicteatime):

hmmm....

OpenStudy (epicteatime):

1-99/7=271? xD

OpenStudy (firekat97):

can we use the numbers 1, 9, 9 and 7 as many times as we'd like?

OpenStudy (epicteatime):

nope xD

OpenStudy (epicteatime):

sadly xD

OpenStudy (firekat97):

damn ahahha

OpenStudy (firekat97):

but can we rearrange stuff to make like 99 and 17 and stuff?

OpenStudy (epicteatime):

99 yes, 17 no xD

OpenStudy (epicteatime):

you have to keep 1997 in that order